OpenStudy (thedj4jc):

Find the quotient. (6x2 - x - 40) ÷ (5 + 2x) 8 - 3x 3x + 8 3x - 8

OpenStudy (freckles):

first can you simplify 6x^2/2x?

OpenStudy (dtan5457):

One way to figure this out is using your answer choices, and foiling it with (5+2x) to see if you get back 6x^2-x-40.

OpenStudy (thedj4jc):

3x^2??? @freckles

OpenStudy (freckles):

well x^2/x isn't x^2

OpenStudy (freckles):

but the 3 part is right

OpenStudy (freckles):

\[\frac{6x^2}{2x}=\frac{6}{2} \frac{x^2}{x} =\frac{6}{2} \frac{x \cdot x}{x}= \frac{6}{2} \frac{x \cdot x}{x \cdot 1} =\frac{6}{2} \cdot \frac{x}{x} \cdot \frac{x}{1}=?\]

OpenStudy (thedj4jc):

I'm not sure...

OpenStudy (freckles):

6/2=3 x/x=?

OpenStudy (freckles):

and x/1=x of course since anything divided by 1 is the anything

OpenStudy (freckles):

but x/x=1 don't you agree?

OpenStudy (freckles):

isn't anything divided by itself 1?

OpenStudy (thedj4jc):

Yes

OpenStudy (freckles):

such as 5/5=1 6/6=1 and so on...

OpenStudy (freckles):

\[\frac{6x^2}{2x}=\frac{6}{2} \frac{x^2}{x} =\frac{6}{2} \frac{x \cdot x}{x}= \frac{6}{2} \frac{x \cdot x}{x \cdot 1} =\frac{6}{2} \cdot \frac{x}{x} \cdot \frac{x}{1}=3(1)x=3x\]

OpenStudy (freckles):

or you could have simplify it using your rules \[\frac{6x^2}{2x}=\frac{6}{2} \frac{x^2}{x}=\frac{6}{2} \frac{x^2}{x^1}=3x^{2-1}=3x^1=3x\]

OpenStudy (freckles):

so this is what we have so far